1. For the ARIMA model below, give values for IE (Yt) and Var (Yt).

Yt = Yt−1 − 0.08Yt−2 + et − 0.02et−1?
Answer #1

For ARIM A mode below evaluate Ecve) V (Y) it = Yt-1 -0.08 YT- tet -0.02(t-1 E[vc) = E(Y2-1) - 0.08 E(Y6-2) + Elet) - 0.02 et